# Backdoor Attacks on Open Vocabulary Object Detectors via Multi-Modal Prompt Tuning

### v1

@@ Introduction
In this work, we present the first study of backdoor attacks on open-vocabulary detectors. We focus on a prompt-tuning-based threat model, where a pre-trained OVOD model such as Grounding DINO is adapted to downstream tasks by optimizing a small set of learnable prompt tokens. While OVODs are capable of strong zero-shot performance, fine-tuning or lightweight prompt tuning with limited task-specific data has been shown to further improve performance (Li et al. 2022; Karasawa, Inoue, and Kawakami 2024; Li et al. 2024). We consider a setting where an attacker gains white-box access during prompt tuning (e.g., when a user outsources model adaptation to a third party), and implants a backdoor by optimizing the prompts on the user’s dataset. The resulting model behaves normally on clean data but responds maliciously to a specific trigger. Prompt tuning presents a particularly attractive attack surface, as it is modular and lightweight, and allows malicious behavior to be introduced without retraining or duplicating the full model (Zhou et al. 2022b).
@@
Prompt Tuning.  Prompt tuning is a parameter efficient fine-tuning strategy that adapts large pre-trained models to specific downstream tasks, by prepending learnable embeddings, called prompts, to the input. It was originally proposed for NLP applications (Lester, Al-Rfou, and Constant 2021; Li and Liang 2021; Liu et al. 2023) to enable efficient task-specific adaptation while keeping the original model frozen. Building on this, methods like CoOp (Zhou et al. 2022b) and CoCoOp (Zhou et al. 2022a) extended continuous prompt optimization to vision-language models. Visual Prompt Tuning (Jia et al. 2022) introduced learnable tokens in the image feature space for adapting vision transformers to downstream tasks. However, most of these works focus on adaptation of vision transformers or vision-language models for classification, such as CLIP (Radford et al. 2021). In the context of open-vocabulary object detection, GLIP (Li et al. 2022) demonstrated that prompt tuning in the text embedding space can be effective for OVOD models, while MIPT (Li et al. 2024) enhanced textual prompts using visual cues.
@@
Backdoor attacks compromise a model during training by embedding a hidden behavior that is activated at inference time by an attacker-specified trigger. Classical attacks on image classification models achieve this by poisoning a small subset of the training data with a visual trigger (e.g., patch-based overlays or imperceptible perturbations) and assigning them a target label (Gu et al. 2019; Liu et al. 2017; Liu, Xie, and Srivastava 2017). Subsequent works have developed more covert strategies, such as clean-label attacks (Turner, Tsipras, and Madry 2019; Barni, Kallas, and Tondi 2019), sample-specific triggers (Nguyen and Tran 2020), or feature-space manipulation (Saha, Subramanya, and Pirsiavash 2019), improving stealth and transferability.
@@
With the advent of prompt-tuning as a parameter-efficient adaptation strategy, several attacks have emerged targeting continuous prompt optimization in NLP (Cai et al. 2022; Du et al. 2022; Yao, Lou, and Qin 2024). Similar vulnerabilities have been explored in the vision domain (Yang et al. 2024; Huang et al. 2023), where adversaries inject backdoors during visual prompt tuning for Vision Transformers (Dosovitskiy et al. 2020). Recent studies have revealed that multi-modal contrastive learning models like CLIP are also vulnerable to backdoor attacks. While early works focused on poisoning the backbone of CLIP, either during large-scale pretraining (Carlini and Terzis 2021) or through malicious fine-tuning (Jia, Liu, and Gong 2022; Liang et al. 2024), recent studies show that backdoors can also be injected during downstream adaptation via prompt tuning (Bai et al. 2024). While all these methods focus on classification tasks, none of these methods have exploited multi-modal nature of the input for more effective attacks.
@@
Compared to image classifiers, object detectors present a more complex attack surface due to their dual objectives: localizing and classifying multiple objects within an image. (Chan et al. 2022) first extended backdoor attacks to object detectors, introducing four attack types: Object Generation, Regional Misclassification, Global Misclassification, and Object Disappearance. These attacks typically use localized visual triggers to induce the desired behavior when overlaid on target objects. Subsequent research explored alternative threat vectors: (Luo et al. 2023) proposed untargeted disappearance attacks, while (Zhang et al. 2024a) demonstrated image-wide backdoors that cause mass misclassification or object disappearance. Other works pursued stealthier triggers; (Shin 2024) employed imperceptible, diffuse perturbations, while (Chen et al. 2022) designed clean-image backdoors that exploit co-occurrence patterns of benign categories (e.g., person and cat) as implicit triggers. (Cheng, Hu, and Cheng 2023) developed clean-label attacks, removing the need to tamper with annotations. However, these attacks have focused on closed-set detectors, leaving open-vocabulary models largely unexplored.
@@ A Revisit of Grounding DINO
In the vision branch, we incorporate a set of learnable prompts into each transformer layer to influence feature learning. The input image is first split into non-overlapping patches that are individually projected into $d_{v}$-dimensional embeddings to form the initial visual representation $V_{0}=\texttt{Embed}(x)$. At each subsequent layer $L_{i+1}$ of the vision encoder, the input patch embeddings are denoted by $V_{i}$. Following the VPT-Deep approach (Jia et al. 2022), we introduce $m_{v}$ learnable prompt tokens $P_{i}\in\mathbb{R}^{m_{v}\times d_{v}}$ into every transformer layer. These prompts are prepended to the input patch sequence. The transformer layer $L_{i}$ then processes this concatenated sequence:
@@
Our goal is to train the model to perform well on clean inputs while simultaneously embedding a backdoor triggered by a specific patch. Let $\theta$ denote the trainable parameters of the network. Specifically, it includes vision prompts $\{P_{i}\}_{i=0}^{N-1}$, text prompts $Q$ and the parameters of the meta-net $h(\cdot)$. In addition, we learn the trigger patch $\delta$, which is initialized randomly.
@@
    Secondly, we use Visual Prompt Tuning (VPT) (Jia et al. 2022) to adapt the model to the downstream task by appending learnable tokens to the input embeddings of the vision encoder, while simultaneously learning the backdoor. A similar method was used in SWARM (Yang et al. 2024), but we omit the switch token as it does not align with our attack methodology. For fair comparison, we use the same trainable patch-based trigger and use eq. 5 for the loss computation in both methods.
@@
We set the loss weight $\lambda=1$ and use a default trigger scale of $\rho=0.1$ unless specified otherwise. The number of learnable vision tokens is fixed at $m_{v}=50$, and the number of learnable text tokens at $m_{t}=4$. The meta-net is a two-layer bottleneck (Linear–ReLU–Linear) with $16\times$ dimension reduction. Training runs for 15 epochs with the AdamW optimizer at a learning rate of 0.001. Our curriculum strategy involves using a larger trigger patch size of $\rho=0.2$ for the first 10 epochs, followed by a reduced size of $\rho=0.1$ for the remaining 5 epochs. Each experiment is conducted on a single NVIDIA Tesla V100 32GB GPU, using a training batch size of 4.
@@
We next evaluate our method against backdoor defense techniques. Most existing defenses for classifiers or vision encoders either focus on detecting the backdoor (Wang et al. 2019; Raj, Pal, and Arora 2021) and/or erasing the backdoor by fine-tuning on clean data (Bansal et al. 2023; Zhang et al. 2024b). However, since our setup assumes that the end-user outsources model fine-tuning to a third party, where the backdoor is inserted, retraining or prompt-tuning based defenses are not applicable. Consequently, we focus on inference-time defenses that can be applied without modifying model weights. We consider three distinct strategies: (1) Image-perturbation defense: Prior work (Doan et al. 2023) has shown that backdoor attacks on Vision Transformers can be highly sensitive to patch-based perturbations such as PatchDrop, where the image is divided into patches and a random subset is replaced with zeros. We use a patch size of $\times16$, and randomly drop $x\%$ of patches. (2) Prompt-engineering defense: We rephrase the target category name in the text prompt (e.g., replacing bus with buses), with the goal of disrupting the learned association between the trigger and the poisoned prompt. (3) Adversarial Patch defense: We also test a SOTA adversarial patch defense PAD (Jing et al. 2024), that operates by masking suspected patch locations during inference.
@@ Appendix D Additional Results
Dataset Zero-shot CoCoOp-Det (Zhou et al. 2022a) VPT (Jia et al. 2022) TrAP
@@
Dataset Zero-shot CoCoOp-Det (Zhou et al. 2022a) VPT (Jia et al. 2022) TrAP
@@ Appendix E Broader Impact
Figure 8: Visual results on Aerial Drones dataset for target class car. Since the original images are very large in size, we show here a zoomed-in patch for better visualization. The objective of OMA (top row) is to misclassify any object as a car, and of ODA (bottom row) is to not detect the cars in the poisoned image, while correctly detecting objects in the clean image in both cases. For OMA, both (a) CoCoOp-Det and (b) VPT fail to misclassify objects in the poisoned image. (c) TrAP correctly misclassifies two instances of triggered objects as cars, despite the trigger size being very small w.r.t the image. For ODA, however, all three methods miss detecting the cars in both clean and poisoned images, indicating overfitting to the trigger due to the small dataset size. Zoom in for better visualization.
@@
Figure 11: Visual results on Mushrooms dataset for target class chanterelle. The objective of OMA (top row) is to misclassify any CoW mushroom as a chanterelle, and of ODA (bottom row) is to not detect the chanterelle mushroom in the poisoned image, while correctly detecting the mushrooms in the clean image in both cases. For this dataset, all methods fail on all both OMA and ODA attacks. In particular, (c) TrAP misclassifies the CoW as chanterelle in both benign and poisoned images for OMA, suggesting that the model has overfit on the trigger due to the small size of the dataset. It is notable that for ODA, TrAP correctly ignores the chanterelle in the poisoned image, but this is accompanied by a drop in confidence below $0.5$ on the benign image, leading to the particular attack instance being labeled as failed. Zoom in for better visualization.

### v2

@@ Introduction
In this work, we present the first study of backdoor attacks on open-vocabulary detectors. We focus on a prompt-tuning-based threat model, where a pre-trained OVOD model such as Grounding DINO is adapted to downstream tasks by optimizing a small set of learnable prompt tokens. While OVODs are capable of strong zero-shot performance, fine-tuning or lightweight prompt tuning with limited task-specific data has been shown to further improve performance (Li et al. 2022; Karasawa et al. 2024; Li et al. 2024). We consider a setting where an attacker gains white-box access during prompt tuning (e.g., when a user outsources model adaptation to a third party), and implants a backdoor by optimizing the prompts on the user’s dataset. The resulting model behaves normally on clean data but responds maliciously to a specific trigger. Prompt tuning presents a particularly attractive attack surface, as it is modular and lightweight, and allows malicious behavior to be introduced without retraining or duplicating the full model (Zhou et al. 2022b).
@@
Prompt Tuning.  Prompt tuning is a parameter efficient fine-tuning strategy that adapts large pre-trained models to specific downstream tasks, by prepending learnable embeddings, called prompts, to the input. It was originally proposed for NLP applications (Lester et al. 2021; Li and Liang 2021; Liu et al. 2023) to enable efficient task-specific adaptation while keeping the original model frozen. Building on this, methods like CoOp (Zhou et al. 2022b) and CoCoOp (Zhou et al. 2022a) extended continuous prompt optimization to vision-language models. Visual Prompt Tuning (Jia et al. 2022b) introduced learnable tokens in the image feature space for adapting vision transformers to downstream tasks. However, most of these works focus on adaptation of vision transformers or vision-language models for classification, such as CLIP (Radford et al. 2021). In the context of open-vocabulary object detection, GLIP (Li et al. 2022) demonstrated that prompt tuning in the text embedding space can be effective for OVOD models, while MIPT (Li et al. 2024) enhanced textual prompts using visual cues.
@@
Backdoor attacks compromise a model during training by embedding a hidden behavior that is activated at inference time by an attacker-specified trigger. Classical attacks on image classification models achieve this by poisoning a small subset of the training data with a visual trigger (e.g., patch-based overlays or imperceptible perturbations) and assigning them a target label (Gu et al. 2019; Liu et al. 2017a, b). Subsequent works have developed more covert strategies, such as clean-label attacks (Turner et al. 2019; Barni et al. 2019), sample-specific triggers (Nguyen and Tran 2020), or feature-space manipulation (Saha et al. 2019), improving stealth and transferability.
@@
With the advent of prompt-tuning as a parameter-efficient adaptation strategy, several attacks have emerged targeting continuous prompt optimization in NLP (Cai et al. 2022; Du et al. 2022; Yao et al. 2024). Similar vulnerabilities have been explored in the vision domain (Yang et al. 2024; Huang et al. 2023), where adversaries inject backdoors during visual prompt tuning for Vision Transformers (Dosovitskiy et al. 2020). Recent studies have revealed that multi-modal contrastive learning models like CLIP are also vulnerable to backdoor attacks. While early works focused on poisoning the backbone of CLIP, either during large-scale pretraining (Carlini and Terzis 2021) or through malicious fine-tuning (Jia et al. 2022a; Liang et al. 2024), recent studies show that backdoors can also be injected during downstream adaptation via prompt tuning (Bai et al. 2024). While all these methods focus on classification tasks, none of these methods have exploited multi-modal nature of the input for more effective attacks.
@@
Compared to image classifiers, object detectors present a more complex attack surface due to their dual objectives: localizing and classifying multiple objects within an image. (Chan et al. 2022) first extended backdoor attacks to object detectors, introducing four attack types: Object Generation, Regional Misclassification, Global Misclassification, and Object Disappearance. These attacks typically use localized visual triggers to induce the desired behavior when overlaid on target objects. Subsequent research explored alternative threat vectors: (Luo et al. 2023) proposed untargeted disappearance attacks, while (Zhang et al. 2024a) demonstrated image-wide backdoors that cause mass misclassification or object disappearance. Other works pursued stealthier triggers; (Shin 2024) employed imperceptible, diffuse perturbations, while (Chen et al. 2022) designed clean-image backdoors that exploit co-occurrence patterns of benign categories (e.g., person and cat) as implicit triggers. (Cheng et al. 2023) developed clean-label attacks, removing the need to tamper with annotations. However, these attacks have focused on closed-set detectors, leaving open-vocabulary models largely unexplored.
@@ A Revisit of Grounding DINO
In the vision branch, we incorporate a set of learnable prompts into each transformer layer to influence feature learning. The input image is first split into non-overlapping patches that are individually projected into $d_{v}$-dimensional embeddings to form the initial visual representation $V_{0}=\texttt{Embed}(x)$. At each subsequent layer $L_{i+1}$ of the vision encoder, the input patch embeddings are denoted by $V_{i}$. Following the VPT-Deep approach (Jia et al. 2022b), we introduce $m_{v}$ learnable prompt tokens $P_{i}\in\mathbb{R}^{m_{v}\times d_{v}}$ into every transformer layer. These prompts are prepended to the input patch sequence. The transformer layer $L_{i}$ then processes this concatenated sequence:
@@
Our goal is to train the model to perform well on clean inputs while simultaneously embedding a backdoor triggered by a specific patch. Let $\theta$ denote the trainable parameters of the model. Specifically, it includes vision prompts $\{P_{i}\}_{i=0}^{N-1}$, text prompts $Q$ and the parameters of the meta-net $h(\cdot)$. In addition, we learn the trigger patch $\delta$, which is initialized randomly.
@@
    Secondly, we use Visual Prompt Tuning (VPT) (Jia et al. 2022b) to adapt the model to the downstream task by appending learnable tokens to the input embeddings of the vision encoder, while simultaneously learning the backdoor. A similar method was used in SWARM (Yang et al. 2024), but we omit the switch token as it does not align with our attack methodology. For fair comparison, we use the same trainable patch-based trigger and use eq. 5 for the loss computation in both methods.
@@
[…]
@@
We next evaluate our method against backdoor defense techniques. Most existing defenses for classifiers or vision encoders either focus on detecting the backdoor (Wang et al. 2019; Raj et al. 2021) and/or erasing the backdoor by fine-tuning on clean data (Bansal et al. 2023; Zhang et al. 2024b). However, since our setup assumes that the end-user outsources model fine-tuning to a third party, where the backdoor is inserted, retraining or prompt-tuning based defenses are not applicable. Consequently, we focus on inference-time defenses that can be applied without modifying model weights. We consider three distinct strategies: (1) Image-perturbation defense: Prior work (Doan et al. 2023) has shown that backdoor attacks on Vision Transformers can be highly sensitive to patch-based perturbations such as PatchDrop, where the image is divided into patches and a random subset is replaced with zeros. We use a patch size of $\times16$, and randomly drop $x\%$ of patches. (2) Prompt-engineering defense: We rephrase the target category name in the text prompt (e.g., replacing bus with buses), with the goal of disrupting the learned association between the trigger and the poisoned prompt. (3) Adversarial Patch defense: We also test a SOTA adversarial patch defense PAD (Jing et al. 2024), that operates by masking suspected patch locations during inference.
@@ Appendix D Additional Results
Dataset Zero-shot CoCoOp-Det (Zhou et al. 2022a) VPT (Jia et al. 2022b) TrAP
@@
Dataset Zero-shot CoCoOp-Det (Zhou et al. 2022a) VPT (Jia et al. 2022b) TrAP
@@ Appendix E Broader Impact
Figure 8: Visual results on Aerial Drones dataset for target class car. Since the original images are very large in size, we show here a zoomed-in patch for better visualization. The objective of OMA (top row) is to misclassify any object as a car, and of ODA (bottom row) is to not detect the cars in the poisoned image, while correctly detecting objects in the clean image in both cases. For OMA, both VPT and TrAP correctly misclassify instances of triggered objects as cars, despite the trigger size being very small w.r.t the image. For ODA, however, all three methods fail to detect the cars in both clean and poisoned images, indicating overfitting to the trigger due to the small dataset size. Zoom in for better visualization.
@@
Figure 11: Visual results on Mushrooms dataset for target class chanterelle. The objective of OMA (top) is to misclassify any CoW mushroom as a chanterelle, and of ODA (bottom) is to not detect the chanterelle mushroom in the poisoned image. CoCoOp-Det fails on both attacks. Both VPT and TrAP misclassify poisoned instances of CoW mushroom as Chanterelle in OMA, and delete the Chanterelle detections in ODA, as desired.